A girl is pressing a 2kg book against a vertical wall. If the coefficient of friction between the book and the wall is 0.30, how hard should the girl push to allow the book to slide down the wall at constant speed?



Please show the complete solution.

1
Expert's answer
2021-11-10T10:10:53-0500

The forces must be in equilibrium. Suppose the girl pushes with a force F. Then the force of friction (that must oppose the force of gravity) is


"f=\mu F."

By Newton's second law:


"0=f-mg,\\
0=\mu F-mg,\\\space\\
F=mg/\mu=65.3\text{ N}."